OpenStudy (nevermind_justschool):

@CoDsTeR9000 Ask your questions and I will help!

OpenStudy (nevermind_justschool):

@omgitsjc

OpenStudy (gnrfan):

As one of the new roller coaster engineers, you have been tasked with developing a roller coaster that will intertwine with existing Oakville Lake Amusement Park structures. For one of the more thrilling sections, the roller coaster will dive down in-between buildings, plummet underground, pop back up, and coast over a hill before shooting back underground. There must be three distinct points where the roller coaster crosses the x–axis. Precise measurements and attention to detail are very important. coordinate plane with buildings blocking off x–intercepts of negative 11, negative 10, 0, 1, 2, 3, 9, and 10 First, here is the existing map of current structures. It is important that the roller coaster does not go through the foundation of any of these structures. 1st point: ___6___ 2nd point:___-2___ 3rd point: ___-7___ Using the points above as zeros, construct the polynomial function, f(x), that will be the path of your roller coaster. Show all of your work. Using both fundamental Theorem and Descartes` rule of signs, prove to the construction foreman that your funtion matches your graph. Use complete sentences. Solve for the y–intercept for your function, f(x), and then construct a rough graph of your roller coaster. If your y–intercept is off the graph, give the coordinates of the y–intercept.

OpenStudy (gnrfan):

can u help with this?

OpenStudy (acxbox22):

please post your question separately

OpenStudy (anonymous):

ok im here

OpenStudy (nevermind_justschool):

Finally

OpenStudy (anonymous):

sorry (sarcastic voice)

OpenStudy (anonymous):

so what?

OpenStudy (nevermind_justschool):

Got a question?

OpenStudy (anonymous):

yea just a sec...... here: Solve for x: -2(x + 3) = -2(x + 1) - 4.

OpenStudy (nevermind_justschool):

Alright! Using what you know about solving equations, what do you do first??

OpenStudy (anonymous):

i did it over already and they equal out but idk if i did it right...

OpenStudy (anonymous):

i distributed what was outside of the par. to whats inside the par.

OpenStudy (anonymous):

par.= parentheses

OpenStudy (nevermind_justschool):

Hey, what was your answer?

OpenStudy (anonymous):

-2x+(-6)=-2x+(-6)

OpenStudy (anonymous):

back

OpenStudy (nevermind_justschool):

Okay Just a sec. I am doing it on paper

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answers that i can have for it are A: 2 B: 3 C: All real numbers D: No solution

OpenStudy (anonymous):

i want to say either all real numbers or no solution

OpenStudy (nevermind_justschool):

C. I think, I am not sure though...Sorry I couldn't be of more assistance

OpenStudy (anonymous):

its fine cuz i was pullin towards that answer anyways so yea thanks! but stay cuz i have some more to ask about

OpenStudy (anonymous):

i will medal u though and fan so yea!

OpenStudy (nevermind_justschool):

Alright-y then! Thanks :)

OpenStudy (anonymous):

ok so i have one more and then i might have another one so yea can u stay a little longer?

OpenStudy (nevermind_justschool):

I guess

OpenStudy (anonymous):

ok so i have a word prob. its kinda tricky so yea here it is Eddie has saved up $45 to purchase a new camera from the local store. The sales tax in his county is 7% of the sticker price. Write an equation and solve it to determine the value of the highest priced camera Eddie can purchase with his $45, including the sales tax. Round your answer to the nearest penny.

OpenStudy (anonymous):

and one personal question.... do u live in florida? not ment to be creepy...just saying

OpenStudy (nevermind_justschool):

Nope :) I don't live in FL :)

OpenStudy (anonymous):

dang well ok then :)

OpenStudy (nevermind_justschool):

Alright, what is your answer for the problem?

OpenStudy (anonymous):

well ive tried on it and got nuthin.... just nuthin cuz im stuck on tryin to make an equation but cant seem to make one... :(

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

well the answers i can have are A: x - 0.07x = 45; $48.39 B: 0.7x = 45; x = $64.29 C: 7x = 45; x = $6.43 D: x + 0.07x = 45; x = $42.06

OpenStudy (anonymous):

not giving u a guilt trip in the msg ok?

OpenStudy (nevermind_justschool):

Alright ;) Working the problem out Be done in a sec

OpenStudy (anonymous):

ok ive been working on it too while u were gone and i want to pull towards either A or C

OpenStudy (nevermind_justschool):

I think C.

OpenStudy (anonymous):

yea.... probably. :) thnx now i might have another one so hold up just a sec but if not then yea so brb :)

OpenStudy (nevermind_justschool):

Okay

OpenStudy (anonymous):

yea its brobably easy to u cuz ita a fraction/divition kind of problem just a sec

OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x. like x+20/2=3x

OpenStudy (nevermind_justschool):

I don't understand... What is the question?

OpenStudy (anonymous):